# Geometric Signatures of Compositionality Across a Language Model’s Lifetime

### v2

@@ 3.3 Measuring feature complexity via dimensionality estimation
We report the nonlinear $I_{d}$ using the TwoNN estimator of Facco et al., 2017. We choose TwoNN as opposed to other measures of nonlinear dimensionality for several reasons. First, it is highly correlated to other state-of-the-art estimators, such as the Maximum Likelihood Estimator (MLE) of Levina & Bickel (2004) , for both synthetic point cloud benchmarks (Facco et al., 2017) and LM representations (Cheng et al., 2023). Second, it relies on minimal assumptions of local uniformity up to the second nearest neighbor of a point, in contrast to other estimators that impose stricter assumptions, for instance, global uniformity (Albergante et al., 2019). Third, TwoNN and correlated estimators enjoy precedence in related manifold estimation literature (Cheng et al., 2023; Pope et al., 2021; Chen et al., 2024; Tulchinskii et al., 2023; Ansuini et al., 2019). In addition to TwoNN in the main text, we also test MLE in Appendix C, confirming they are highly correlated.
The TwoNN estimator works as follows. Points on the underlying manifold are assumed to follow a locally homogeneous Poisson point process. Here, local refers to the neighborhood about each point $x$ encompassing $x$’s first and second nearest neighbors. Let $r_{k}^{(i)}$ be the Euclidean distance between $x_{i}$ and its $k$th nearest neighbor. Then, under the mentioned assumptions, the distance ratios $\mu_{i}:=r_{2}^{(i)}/r_{1}^{(i)}\in[1,\infty)$ follow the cumulative distribution function $F(\mu)=(1-\mu^{-I_{d}})\mathbf{1}[\mu\geq 1]$. This yields an estimator for the ID, $I_{d}=-\log(1-F(\mu))/\log\mu$. Finally, given representations $\{x_{i}^{(j)}\}_{i=1}^{N}$ for LM layer $j$, $I_{d}^{(j)}$ is numerically fit via maximum likelihood estimation over all datapoints.
@@
To estimate the linear effective dimension $d$, we use Principal Component Analysis (PCA) (Jolliffe, 1986) with a variance cutoff of $99\%$. We compared to the Participation Ratio (PR) (Gao et al., 2017), a linear dimensionality measure often used in the computational neuroscience literature (cf. Chung et al. (2018); Recanatesi et al. (2019)), finding it to produce uninterpretable results, see Appendix C. For this reason, we focus on PCA in the main text.
Figure 2: Mean dimensionality over model size. Mean nonlinear $I_{d}$ (left) and linear $d$ (right) over layers is shown for increasing LM hidden dimension. While nonlinear $I_{d}$ does not depend on hidden dimension $D$ (flat lines), PCA $d$ scales linearly in $D$. Curves are averaged over 5 data splits, $\pm$ 1 SD.
@@ 4 Results
We find that representational dimensionality reflects compositionality in ways that are predictable over pre-training and model scale. First, we show that language models represent linguistic data on low-dimensional, nonlinear manifolds, but in high-dimensional linear subspaces that scale linearly with the hidden dimension. Then, we show that, over training, geometric feature complexity is informative of an LM’s linguistic competence, such that both exhibit a nontrivial phase transition that marks emergence of syntactic and semantic abilities. Finally, we show that representational dimensionality predictably reflects the degree of compositionality, both in terms of combinatorial complexity and sequence-level semantics and analyze its evolution over training. For brevity, we focus on model sizes 410m, 1.4b, and 6.9b in the main text, with full results in the appendix.
@@ 4.1 Nonlinear and linear feature complexity scale differently with model size
Like in previous work (Cai et al., 2021; Valeriani et al., 2023; Cheng et al., 2024), we confirm that input data are represented on a nonlinear manifold with orders-of-magnitude lower dimension than the embedding dimension. In particular, for both the controlled dataset, see Figure 2, and for The Pile, see Figure H.1, we find that $I_{d}\sim O(10)$ while linear $d,D\sim O(10^{3})$.
@@
These results highlight key differences in how linear and nonlinear dimensions are recruited: LMs globally distribute representations to occupy $d\propto D$ dimensions of the space, but their shape is locally constrained to a low-dimensional ($I_{d}$) manifold. Robustness of $I_{d}$ to scaling the hidden dimension suggests that LMs, once sufficiently performant, recover the degrees of freedom underlying the data.
@@
Figure 3: ID tracks task performance. Top: Layerwise $I_{d}$ development of Pythia-410m, 1.4b, and 6.9b over pre-training. The phase transition of ID around checkpoint $10^{3}$ is persistent across the model sizes. Bottom: Zero-shot task performance of various LM evaluation tasks of the same models across pre-training. Also around checkpoint $10^{3}$, linguistic competence measured by task performance starts to increase for all models.
@@
4.2 Evolution of representational geometry tracks emergent linguistic abilities over training
We just saw how dimensionality scales with size, and now we investigate its change over time. We find that feature complexity is highly related to the LM’s linguistic capabilities, assessed using the eval-harness benchmark performance, over training. Figure 3 shows the evolution of $I_{d}$ on the $k=1$ dataset (top), where each curve is one layer, with the evolution of LM performance on the benchmark tasks (bottom), where each curve plots performance on an individual task.
@@ 4.2 Evolution of representational geometry tracks emergent linguistic abilities over training
We observe in Figure 3 that, for all models, the evolution of representational dimensionality closely tracks a sudden transition in LM task performance. In Figure 3 top, we first observe that $I_{d}$ decreases sharply before checkpoint $10^{3}$ and then re-distributes. At the same time, task performance rapidly improves after the steep decrease in $I_{d}$ (Figure 3 bottom). Feature complexity evolution on The Pile is shown in Figure H.2, and exhibits a similar transition to that reported in Figure 3. Further, the existence of the phase transition in representational geometry $t\approx 10^{3}$ is robust to the dimensionality measure and whether the data are shuffled, see Figure G.4. Our results resonate with Chen et al. (2024), who observed in BERT models a similar two-part $I_{d}$ transition on the training corpus; they showed that the two extrema corresponding to the dip and uptick in $I_{d}$ temporally coincided with the onset of higher-order linguistic capabilities. Together, results show that representational complexity can signify whether and when LMs learn linguistic structure. Crucially, we show that the phase transition exists for inputs beyond in-distribution data, which was the subject of (Chen et al., 2024), and, furthermore, beyond grammatical data (Figure G.4) as a more general property of LM processing.
4.3 Representational complexity reflects input compositionality
@@
We just established that feature complexity is informative of when models gain complex linguistic capabilities that, by definition, require compositional understanding. Now, we establish our key result, which is that feature complexity encodes input compositionality, both when considering formal compositionality, or data combinatorial complexity, as well as meaning compositionality, or sentence-level semantics. We first show that this holds for fully-trained models that have reached final linguistic competency. Then, using evidence from the training phase of the model, we show that the correspondence between feature complexity and input compositionality is present first as an inductive bias of the model that encodes formal complexity; but then, that it persists due to learned syntactic and semantic features that encode meaning complexity. Lastly, we further develop the coding differences between $d$ and $I_{d}$, confirming an existing hypothesis in the literature (Recanatesi et al., 2021) that they respectively encode formal and semantic complexity of inputs.
@@
Figure 4: Dimensionality over layers. Nonlinear $I_{d}$ (top) and linear $d$ (bottom) over layers are shown for sizes 410m, 1.4b, and 6.9b (left to right). Each color corresponds to a coupling length $k\in 1\cdots 4$. Solid curves denote sane sequences, and dotted curves denote shuffled sequences. For all models, lower $k$ results in higher $I_{d}$ and $d$ for both normal and shuffled settings. For all models, shuffling results in lower $I_{d}$ but higher $d$. Curves are averaged over 5 random data splits, shown with ±$\pm 1$ SD (shaded); SDs across random data splits tended to be very small.
Data combinatorial complexity
@@
On fully-trained models, representational dimensionality preserves relative dataset compositionality. Figure 4 shows for fully-trained Pythia 410m, 1.4b, and 6.9b that $I_{d}$ and $d$ increase with the degree of formal compositionality within both sane (solid) and shuffled (dashed) settings: the highest curves (blue) correspond to the $k=1$ dataset, or 12 degrees of freedom, and the lowest (red) denote the $k=4$ dataset, or 3 degrees of freedom. The relative order of feature complexity, moreover, holds for all layers, seen by non-overlapping solid curves in Figure 4.
@@
Grammaticality is not a precondition for representational dimensionality to reflect data combinatorial complexity: in Figure 4 (top), dashed curves corresponding to shuffled text are also ordered $k=1\cdots 4$ top to bottom. While the relative order of formal complexity is preserved in the LM’s feature complexity for both grammatical and agrammatical datasets, the separation is greater for grammatical text (solid curves). We hypothesize that this is due to shuffled text being out-of-distribution, such that the model cannot integrate the sequences’ meaning, but nevertheless preserve surface-level complexity in its representations. This tendency holds for pre-trained models of all sizes, see Figure G.1.
@@
The relationship between dimensionality and data combinatorial complexity, controlled by $k$, for sane text is not an emergent feature over training. In Figure 5 (left), the inverse relationship between $k$ and both $I_{d}$ and $d$ is present throughout training. But, the reason for this relation differs at the start and end: in shuffled text, where sequence-level semantics are not present, the relationship between $k$ and dimensionality is salient at the beginning and greatly diminishes by the end, whereas in sane text it remains salient throughout training. Together, these demonstrate an inductive bias of the initialized LM architecture to preserve input complexity in its representations. Then, over training, differences in dimensionality may be increasingly explained by features beyond the surface distribution of inputs. We claim that these features are semantic, providing evidence towards the claim in what follows.
@@
Figure 5: Training dynamics of dimensionality. The top row shows results for TwoNN $I_{d}$ and the bottom row shows the PCA $d$. (Left): Layerwise $I_{d}$ at different timepoints of training for sane vs. shuffled examples with different coupling $k$ (6.9b model). $I_{d}$ difference of shuffled examples with varying $k$ diminishes as the training persists. All curves are shown with ±$\pm 1$ SD (the SDs were very small). (Right): $\Delta I_{d}$ between $k=1,k=4$ across training for various model sizes (different colors).
Compositional semantics
@@
Shuffling sequences destroys their meaning, removing dataset complexity attributed to sentence-level semantics. Figure 4 shows, for fully-trained models, the dimensionality over layers for sane and shuffled inputs. Here, nonlinear and linear dimensionalities show opposing trends: $I_{d}$ for shuffled text collapses to a low range, while $d$ increases, seen by the dashed curves in each plot compared to solid curves. This tendency holds across model sizes, see Figure G.1.
We refer to the phenomenon in which shuffling destroys sequence-level semantics and $I_{d}$, also attested for sequences in Pythia’s training corpus (Cheng et al., 2024), as shuffling feature collapse. Evidence from the training dynamics of the LM further suggests that this feature collapse is due to semantics. We saw in Section 4.2 that training step $t=10^{3}$ approximately marked a phase transition after which the LM’s linguistic competencies sharply rose. Crucially, the epoch $t=10^{3}$ preceding the sharp increase in linguistic capabilities is also the first to exhibit shuffling feature collapse. Figure 5 (right) shows the $\Delta I_{d}$ between the $k=1$ and $k=4$ dataset for several model sizes, across training (x-axis). Shuffling feature collapse, given by low $\Delta I_{d}$, occurs around $t=10^{3}$ for all models. On the other hand, $\Delta I_{d}$ for sane text stabilizes to around $\sim 25$ for different model sizes. This transition does not occur for linear $d$, see Figure 5 (right, bottom). This suggests that shuffling feature collapse for $I_{d}$ is symptomatic of when the LM learns to extract meaningful semantic features.
We interpret shuffling feature collapse using an argument from Recanatesi et al. (2021), who propose that predictive coding requires the model to satisfy two objectives: to encode the vast “space of inputs and outputs", exerting upward pressure on representational complexity, and at the same time, to extract latent features to support prediction, exerting a downward pressure on complexity. Recanatesi et al. (2021) suggest that the first pressure expands the linear representation space $\mathbb{R}^{d}$, while the second compresses representations to a $I_{d}$-dimensional manifold. Indeed, in our setting, shuffling words increases input complexity, thus increasing $d$. But, shuffling destroys sequence semantics, exerting a downward pressure on $I_{d}$. Recanatesi et al. (2021)’s interpretation of linear dimensions as encoding the input space corresponds to what we have been calling formal compositionality; conversely, what they refer to as latent semantic features, encoded nonlinearly, is aligned with our meaning compositionality. We now investigate this form-meaning coding dichotomy in more detail.
Form-meaning dichotomy in representation learning
We proposed in line with Recanatesi et al. (2021) that linear $d$ captures surface-level variation, while $I_{d}$ primarily encodes semantic variation. We have shown the latter in the previous section: $I_{d}$ decreases in the absence of compositional semantics while $d$ does not, suggesting that $I_{d}$, not $d$, encodes sequence-level meaning complexity. If this hypothesis holds, we need to show that linear $d$, and not $I_{d}$, encodes form compositionality.
Form compositionality is quantified by the gzip-compressed size of each dataset. Spearman correlations between gzip (kilobytes), and dimensionality are shown in Table 1. Consistently across model sizes, the average layerwise $I_{d}$ is not correlated to gzip size, while the average layerwise $d$ is highly correlated to gzip size; we discuss the outlier 160m in Appendix I. The high correlation between $d$ and gzip size is, moreover, surprisingly consistent across layers, see Figure I.1, and already present as an inductive bias of the initialized model (see Figure I.2, Appendix I for training dynamics discussion), while the correlation to $I_{d}$ is highly variable and seldom significant for all of training. This suggests that form complexity, already present in the inputs to the LM, is superficially preserved, while meaning complexity is instead constructed, over layers and over training.
@@
Table 1: Spearman correlations between dimensionality and estimated Kolmogorov complexity. The Spearman correlation $\rho$ between the gzipped dataset size (KB) and representational dimensionality (rows), averaged over layers, is shown for all tested model sizes (columns). Values marked with a * are significant with a p-value threshold of 0.05. Values marked with a † are significant with a p-value threshold of 0.1. Generally across models, average-layer $I_{d}$ is not correlated to the estimated Kolmogorov complexity, or formal compositionality, of datasets. Average-layer linear $d$ is consistently highly positively correlated to the estimated Kolmogorov complexity, except one outlier (160m).
Spearman $\rho$ 14m 70m 160m 410m 1.4b 6.9b 12b
$I_{d}$0.20 -0.06 -0.20 -0.05 0.04 0.01 0.05
$d$ 0.90* 0.47 -0.50 0.96* 0.96* 0.92* 0.86*

### v3

@@ 3.3 Measuring feature complexity
We report the nonlinear $I_{d}$ using the TwoNN estimator of Facco et al., 2017. We choose TwoNN for several reasons. First, it is highly correlated to other state-of-the-art estimators for both synthetic point cloud benchmarks (Facco et al., 2017) and LM representations (Cheng et al., 2023). Second, it relies on minimal assumptions of local uniformity up to the second nearest neighbor of a point, in contrast to other estimators that impose stricter assumptions, for instance, global uniformity (Albergante et al., 2019). Third, TwoNN and correlated estimators are commonplace in related manifold estimation literature (Cheng et al., 2023; Pope et al., 2021; Chen et al., 2024; Tulchinskii et al., 2023; Ansuini et al., 2019). We focus on TwoNN in the main text, but also test another $I_{d}$ estimator, Levina and Bickel, 2004’s MLE, in Appendix D, confirming it is highly correlated.
The TwoNN estimator works as follows. Points on the underlying manifold are assumed to follow a locally homogeneous Poisson point process. Here, local refers to the neighborhood about each point $x$, up to $x$’s second nearest neighbor. Let $r_{k}^{(i)}$ be the Euclidean distance between $x_{i}$ and its $k$th nearest neighbor. Then, under the stated assumptions, distance ratios $\mu_{i}\triangleq r_{2}^{(i)}/r_{1}^{(i)}\in[1,\infty)$ follow the cumulative distribution function $F(\mu)=(1-\mu^{-I_{d}})\mathbf{1}[\mu\geq 1]$. This yields the estimator $I_{d}=-\log(1-F(\mu))/\log\mu$. Lastly, given representations $\{x_{i}^{(j)}\}_{i=1}^{N}$ for LM layer $j$, $I_{d}^{(j)}$ is fit via maximum likelihood estimation over all datapoints.
@@
To estimate the linear effective dimension $d$, we use Principal Component Analysis (PCA) (Jolliffe, 1986) with a variance cutoff of $99\%$. For brevity, we focus on PCA in the main text; for comparison to other linear methods (Gao et al., 2017), see Appendix D.
@@ 4 Results
We find representational dimensionality to systematically reflect compositionality over pre-training and model scale. First, we show that LMs represent linguistic data on low-dimensional, nonlinear manifolds, but in high-dimensional linear subspaces that scale linearly with the hidden dimension. Then, we show that, over training, feature complexity tracks an LM’s linguistic competence, where both exhibit a phase transition marking the emergence of syntactic and semantic abilities. Finally, we show that representational dimensionality predictably reflects the degree of compositionality, but that nonlinear $I_{d}$ encodes meaningful compositional complexity while linear $d$ encodes superficial complexity. We focus on Pythia 410m, 1.4b, and 6.9b in the main text, with full results in the Appendix.
@@ 4.1 Nonlinear and linear feature complexity scale differently with model size
We find, as in Cai et al. (2021); Cheng et al. (2023); Valeriani et al. (2023), that LMs represent data on nonlinear manifolds with $I_{d}$ orders-of-magnitude lower than the embedding dimension. For the controlled dataset (Figure 1) and The Pile (Figure H.1), $I_{d}$ is $O(10)$ while linear $d$, $D$ are $O(10^{3})$.
@@
These results highlight key differences in how linear and nonlinear dimensions are recruited: LMs globally distribute representations to occupy $d\propto D$ linear dimensions of the space, but their shape is locally constrained to a low-dimensional ($I_{d}$) manifold. Robustness of $I_{d}$ to model size ($D$) suggests that LMs, once sufficiently performant, recover the degrees of freedom, or compositional complexity, underlying their inputs.
[…]
@@
Figure 2: $I_{d}$ tracks task performance. Top: Layerwise $I_{d}$ over pre-training for Pythia-410m, 1.4b, and 6.9b. The phase transition of $I_{d}$ at $t\approx 10^{3}$ holds across model size. Bottom: Zero-shot task performance across pre-training. Also at $t\approx 10^{3}$, linguistic competence measured by task performance starts to increase for all LMs.
@@
4.2 Feature complexity tracks emergent linguistic abilities over training
Complex linguistic abilities require, by definition, a compositional understanding of language. We use linguistic task performance (Biderman et al., 2023) as a cue for compositional understanding, finding that over training, models’ feature complexity closely tracks the onset of linguistic capabilities.
@@ 4.2 Feature complexity tracks emergent linguistic abilities over training
Figure 2 shows the evolution of $I_{d}$ on the $k$=$1$ dataset (top), where each curve is one layer, with the evolution of LM benchmark performance (bottom), where each curve plots performance on an individual task. For all LMs, the evolution of representational $I_{d}$ closely tracks a sudden transition in task performance. In Figure 2 top, we first observe that $I_{d}$ decreases sharply before checkpoint $10^{3}$ and then redistributes. At the same time $t\approx 10^{3}$, task performance rapidly improves (Figure 2 bottom). Feature complexity evolution on The Pile is shown in Figure H.2, exhibiting a similar transition to that shown in Figure 2. The existence of the phase transition in representational geometry $t\approx 10^{3}$ is robust to the dimensionality measure, and whether the data are shuffled, see Figure G.5. Our results resonate with Chen et al. (2024), who observed in BERT a similar $I_{d}$ transition on the training corpus that coincided with the onset of higher-order linguistic capabilities. Crucially, we show that the phase transition in feature geometry exists for natural language inputs beyond in-distribution data, which was the subject of Chen et al., and, further, beyond grammatical data (Figure G.5) as a more general property of LM processing. Taken together, results show that feature complexity can signify when LMs gain linguistic abilities that require, by definition, compositional understanding, or the ability to construct the meaning of a sentence from the meaning of its parts.
4.3 Feature complexity reflects input compositionality
@@
We just established that feature complexity is informative of when models gain linguistic behaviors that require compositional understanding. Now, we establish our key result, which is that, no matter the model or grammar, feature complexity encodes input compositionality. We first show that this holds for fully-trained models that have reached final linguistic competency. Then, using evidence from the LM’s training phase, we show that the correspondence between feature complexity and input compositionality is present first as an inductive bias of the architecture that encodes superficial input complexity; but then, that it persists due to learned features that encode compositional meaning.
@@
Figure 3: Dimensionality over layers. $I_{d}$ (top) and $d$ (bottom) over layers for Pythia 410m, 1.4b, 6.9b (left to right). Each color is a coupling factor $k\in 1\cdots 4$. Solid lines denote coherent, and dotted curves denote shuffled, text. For all LMs, lower $k$ implies higher $I_{d}$ and $d$ in both coherent and shuffled cases. For all LMs, shuffling collapses the $I_{d}$ but raises $d$. Curves averaged over 5 splits ±$\pm 1$ SD (shaded, SDs very small).
Compositional complexity varying $k$
@@
On all fully-trained models and all datasets, representational dimensionality preserves relative dataset compositionality. Figure 3 shows for fully-trained Pythia 410m, 1.4b, and 6.9b that $I_{d}$ and $d$ increase with the compositionality of coherent inputs (solid curves): the highest curves (blue) correspond to the $k=1$ dataset, or $12$ degrees of freedom, and the lowest (red) denote the $k=4$ dataset, or $3$ degrees of freedom. The relative order of feature complexity, moreover, holds for all layers, seen by non-overlapping solid curves in Figure 3.
@@
Grammaticality is not a precondition for representational dimensionality to reflect data combinatorial complexity: in Figure 3 (top), dashed curves corresponding to shuffled text are also ordered $k=1\cdots 4$ top to bottom. While the relative order of superficial complexity is preserved in the LM’s feature complexity for both grammatical and agrammatical datasets, the separation is greater for grammatical text (solid curves). We hypothesize that this is due to shuffled text being out-of-distribution, such that the model cannot integrate the sequences’ meaning, but nevertheless preserve surface-level complexity in its representations. This tendency holds for pre-trained LMs of all sizes and families (see Figures G.1 and G.2) and for different-length sentences (see Figure J.3).
@@
The relationship between dimensionality and degree of compositionality, controlled by $k$, for coherent text is not an emergent feature over training. In Figure 4 (left), the inverse relationship between $k$ and both $I_{d}$ and $d$ is present throughout training. But, the reason for this relation differs at the start and end: in shuffled text, where phrase-level semantics is ablated, the inverse relationship between $k$ and dimensionality exists at the start and greatly diminishes by the end of training, while in coherent text it is salient throughout. Together, results show an inductive bias of the initialized LM to preserve input complexity in its representations. Then, over training, differences in dimensionality may be increasingly explained by features beyond surface variation of inputs. We claim these features are semantic and provide evidence in what follows.
@@
Figure 4: Training dynamics of dimensionality. (Top) TwoNN $I_{d}$; (Bottom) PCA $d$. (Left): For Pythia-6.9b, layerwise $I_{d}$ for $3$ training checkpoints for coherent vs. shuffled sequences, with different coupling $k$. The $I_{d}$ difference on shuffled data with varying $k$ diminishes as training persists. All curves are shown with ±$\pm 1$ SD (SDs are very small). (Right): $\Delta I_{d}$ between $k=1,k=4$ across training for various model sizes (different colors).
Breaking compositional structure
@@
Shuffling sequences destroys their meaning, removing dataset complexity attributed to sentence-level semantics. Figure 3 shows, for fully-trained models, dimensionality over layers for coherent and shuffled inputs. Here, nonlinear $I_{d}$ and linear $d$ show opposing trends: $I_{d}$ for shuffled text collapses to a low range, while $d$ increases, seen by the dashed curves in each plot compared to solid curves. This trend is robust to sentence length (Figure J.3) and model size (Figure G.1) and family (Figure G.2).
Spearman $\rho$ 14m 70m 160m 410m 1.4b 6.9b 12b Llama Mistral
$I_{d}$0.20 -0.06 -0.20 -0.05 0.04 0.01 0.05 -0.36 0.00
$d$ 0.90* 0.47 -0.50 0.96* 0.96* 0.92* 0.86* 1.0* 1.0*
@@
Table 1: Spearman correlations between dimensionality and estimated Kolmogorov complexity. Spearman correlations $\rho$ between the gzipped dataset size (KB) and representational dimensionality (rows), averaged over layers, for all Pythia sizes (left 7 columns) and Llama and Mistral (right 2 columns). Values with * are significant at a p-value threshold of $0.05$, and † at $0.1$. Across LMs, average-layer $I_{d}$ is not correlated to the estimated KC, or surface complexity, of datasets. Linear $d$ is highly correlated to estimated KC, except the outlier 160m.
We refer to the phenomenon in which shuffling destroys sequence-level semantics and $I_{d}$, also attested for sequences in Pythia’s training corpus (Cheng et al., 2025), as shuffling feature collapse. Evidence from the training dynamics of the LM further suggests that this feature collapse is due to semantics. We saw in Section 4.2 that training step $t=10^{3}$ approximately marked a phase transition after which the LM’s linguistic competencies sharply rose. Crucially, the epoch $t=10^{3}$ preceding the sharp increase in linguistic capabilities is also the first to exhibit shuffling feature collapse. Figure 4 (right) shows the $\Delta I_{d}$ between the $k=1$ and $k=4$ dataset for several model sizes, across training (x-axis). Shuffling feature collapse, given by low $\Delta I_{d}$, occurs around $t=10^{3}$ for all models. On the other hand, $\Delta I_{d}$ for coherent text stabilizes to around $25$ for different model sizes. This transition does not occur for linear $d$, see Figure 4 (right, bottom). This suggests that shuffling feature collapse for $I_{d}$ is symptomatic of when the LM learns to extract meaningful semantic features. We investigate why shuffling feature collapse exists for $I_{d}$, but not for $d$, in the following section. For the emergence of shuffling feature collapse with sequence length, see Appendix J.
5 Form-meaning dichotomy in representation learning
We can interpret shuffling feature collapse using an argument from Recanatesi et al. (2021), who propose that predictive coding requires a model to satisfy two objectives: (1) encode the vast “input and output space", exerting upward pressure on feature complexity, while (2) extracting latent features to support prediction, exerting a downward pressure on complexity. Recanatesi et al. (2021) argue that pressure (1) expands the linear representation space $\mathbb{R}^{d}$, while (2) compresses representations to a $I_{d}$-dimensional manifold. Recall that we quantify a dataset’s superficial complexity by its Kolmogorov complexity. We saw that shuffling words, which increases inputs’ KC, also increases $d$, but that it destroys compositional semantics, collapsing $I_{d}$ (Figure 3). These results suggest a dichotomy in what linear and nonlinear complexity encode, the former capturing superficial complexity (form), and the latter, latent compositional structure (meaning).
We hypothesize that $d$ captures surface variation, while $I_{d}$ encodes meaningful compositional variation. We saw the latter in the previous section: $I_{d}$ collapses to a narrow range in the absence of compositional semantics while $d$ does not, suggesting $I_{d}$, not $d$, encodes compositionality. Now, we show that $d$, not $I_{d}$, encodes superficial complexity. Table 1 shows Spearman correlations $\rho$ between KC (estimated with gzip) and dimensionality. For all LMs, gzip highly correlates to average layerwise $d$ but not to $I_{d}$; we discuss the outlier Pythia-160m in Appendix I. The high $\rho(d,\texttt{gzip})$ is surprisingly consistent across layers (Figures I.2 and I.3) and already present as an inductive bias of the initialized LM (Figure I.5, Appendix I for training discussion). Instead, the correlation to $I_{d}$ is seldom significant for all of training. This suggests surface complexity, already present in the LM’s inputs, is preserved by the architecture, while meaning complexity is instead learned over training.